# Robotic Ultrasound-Guided Femoral Artery Reconstruction of Anatomically-Representative Phantoms

### v1

@@ Abstract
Femoral artery access is essential for numerous clinical procedures, including diagnostic angiography, therapeutic catheterization, and emergency interventions. Despite its critical role, successful vascular access remains challenging due to anatomical variability, overlying adipose tissue, and the need for precise ultrasound (US) guidance. Errors in needle placement can lead to severe complications, restricting the procedure to highly skilled clinicians in controlled hospital settings. While robotic systems have shown promise in addressing these challenges through autonomous scanning and vessel reconstruction, clinical translation remains limited due to reliance on simplified phantom models that fail to capture human anatomical complexity. In this work, we present a method for autonomous robotic US scanning of bifurcated femoral arteries, and validate it on five vascular phantoms created from real patient computed tomography (CT) data. Additionally, we introduce a video-based deep learning US segmentation network tailored for vascular imaging, enabling improved 3D arterial reconstruction. The proposed network achieves a Dice score of 89.21% and an Intersection over Union of 80.54% on a newly developed vascular dataset. The quality of the reconstructed artery centerline is evaluated against ground truth CT data, demonstrating an average $L_{2}$ deviation of 0.91$\pm$0.70 mm, with an average Hausdorff distance of 4.36$\pm$1.11mm. This study is the first to validate an autonomous robotic system for US scanning of the femoral artery on a diverse set of patient-specific phantoms, introducing a more advanced framework for evaluating robotic performance in vascular imaging and intervention.
@@ I Introduction
Femoral arterial access is a fundamental component of interventional medicine, enabling procedures such as diagnostic angiography, therapeutic catheterization, and emergent interventions. Its clinical utilization has expanded over time, with femoral access just for transcatheter aortic valve replacement in the U.S. increasing from 57.1% in 2011 to 95.3% in 2019 [1]. Despite its widespread adoption, femoral access remains one of the most challenging vascular entry sites [2], with a reported 5.9% failure rate in percutaneous aortic procedures [3]. The procedure’s complexity is compounded by anatomical variability, overlying adipose tissue, and the required accuracy in US imaging interpretation. Improper needle placement significantly increases the risk of serious complications, including retroperitoneal hemorrhage, thrombosis, and other vascular sequelae [2, 4]. Given these complexities, femoral arterial access is currently performed only by highly skilled physicians in controlled hospital environments.
Robotic systems can address aforementioned challenges by leveraging their precision and repeatability in US-guided vascular interventions. These systems can enable accurate 3D reconstruction of arterial structures through automated US image segmentation, serving either as fully autonomous platforms, or as clinician-assistive technologies. Furthermore, they present significant potential for deployment in resource-constrained environments where specialized healthcare personnel are unavailable, such as in combat zones, remote locations, or an ambulance. Although robotic US scanning and reconstruction is a mature research topic, most investigations have employed overly simplified phantom models that cannot represent the anatomical complexity encountered in clinical scenarios [5, 6], or relied on single custom-designed phantoms for both system development and validation [7, 8]. Additionally, to the best of the authors’ knowledge, no study has validated robotic femoral artery reconstruction using an independent 3D ground truth derived from a separate imaging modality than that employed by the robotic system itself [9, 10].
@@
In this manuscript, we present an autonomous robotic system for bifurcated femoral artery scanning. The proposed system integrates US image acquisition, arterial segmentation, and 3D centerline reconstruction. In particular, this work’s main contributions are as follows: (1) demonstration of autonomous robotic scanning and 3D reconstruction of bifurcated femoral artery on five vascular phantoms created from real patient computed tomography (CT) data; (2) development of a video-based segmentation network for axial artery segmentation in US video sequences; (3) comprehensive evaluation of the reconstructed bifurcated artery against ground-truth CT data to assess reconstruction accuracy. This study is the first to validate an autonomous robotic system for US scanning of the femoral artery on a diverse set of patient-specific phantoms, introducing a more rigorous framework for evaluating robotic performance in vascular imaging and intervention.
@@ II-1 Vascular Phantoms
Commercial vascular phantoms, while providing good US image quality, assume the form of geometric shapes such as rectangular prisms (CIRS Vascular Access Training Phantom model 72) or semi-cylinders (CAE Vascular Access Phantom) that oversimplify the topology of the scanning surface, and limit the transferability of the system to real human procedures. Though larger models such as the Blue Phantom’s Femoral Access model offer improved anatomical representation, they are often prohibitively expensive. Perhaps most significantly, these phantoms typically derive from a single anatomical model, limiting their representation of human variation, particularly for individuals with higher body mass indices (BMI). Recently, wearable training systems combined with US simulation software, such as the SonoSkin trainer (Simulab, Seattle, WA), garnered more attention. The systems consist of a skin-like cover that can be placed over a mannequin or person, paired with a simulated US probe that triggers predetermined US images at specific anatomical locations. While such systems enable basic examination practice, their reliance on pre-loaded images and lack of orientation-specific feedback limits their utility in developing image acquisition skills, or advancing interpretive technologies. Creating a custom vascular phantom requires developing self-healing materials and durable vessels to withstand repeated needle punctures, a technology largely protected as proprietary by commercial manufacturers. To overcome these challenges, we provide a cost-effective, simple, yet highly effective solution. We develop five novel patient-specific phantoms by modifying the CIRS model $72$ vascular phantom’s surface topology and overall orientation to better represent realistic anatomy, while maintaining the phantom’s established acoustic and mechanical properties.
II-2 Vascular Segmentation in Utrasound
Automatic segmentation of US images acquired during robotic scanning presents significant challenges due to impedance control implementation - while this strategy ensures safe human-robot interaction, it introduces variations in probe contact and pressure. The inherent limitations of US imaging, characterized by poor signal-to-noise ratio and acoustic artifacts, further complicate accurate vascular boundary delineation [11]. Earlier works in US vessel segmentation used geometric and dynamic models, which typically require seed points initialization in the initial frame [12, 13, 14]. To enhance performance, previous works incorporated additional context by using Doppler US with B-mode images [15, 16, 17]. Recently, deep learning has garnered significant attention, with most vessel segmentation studies leveraging UNet-like models that process individual US images [18, 19, 20]. With the advent of transformers for modeling long-range dependencies, several attention mechanisms have been combined with U-Net-like models to encode temporal memory from videos. However, to the best of author’s knowledge, most studies have focused on obstetrics [21], breast sonography [22], and echocardiography [23]. Comparatively limited research has addressed video-based vascular US segmentation, particularly in robotic scanning.
@@ II-3 Robotic Ultrasound Vascular Scanning
Robotic US has emerged as a promising solution to address several drawbacks of hand-held US imaging for trauma [24, 25], abdominal [26, 27] and cardiac [28] procedures. In particular to vascular US imaging, prior research has focused on vessel segmentation, tracking, and localization. Jiang et al. [29] used segmented vascular structures to align a robotic probe perpendicular to the anatomy and later refined spatial positioning using 3D point cloud data of the scanning surface [30]. Chen et al. [31] used color Doppler with B-mode images and analyzed them through deep learning for vascular access. Merouche et al. [7] conducted the automatic tracking of artery and 3D reconstruction by initializing the cross-sectional view of the artery. Ning et al. [32] employed a weakly supervised US vessel segmentation network for probe orientation estimation, and a force-guided reinforcement learning agent to maintain optimal probe angle. However, these systems were validated on either commercial or custom-designed phantoms, which fail to fully capture the anatomical complexity of real human vasculature. While some studies have validated their approaches on animals [9] or human [32] subjects, they were primarily limited to US image acquisition with no independent ground truth for validation. In contrast, our proposed robotic US vascular scanning system is validated on patient-specific phantoms with their corresponding CT scans, a safer and representative testbed for evaluating interventional procedures.
This manuscript is structured as follows: Section III.A presents the methodology and rationale underlying our vascular phantoms development. Section III.B delineates the autonomous robotic US scanning protocol, introducing a novel video-based US segmentation network. Experimental findings from the robotic US images acquisitions are analyzed in Section IV. Finally, Section V synthesizes our conclusions and presents future research directions.
@@ III-A Patient-Specific Vascular Phantoms
The commercial CIRS 72 model was selected as a base vascular phantom, which contains four vessel structures: two vessels of different diameters on the sides, and a central bifurcating vessel that serves as our experimental target. To reduce detection artifacts, the blood-mimicking fluid was drained from both sides of the vessels. We utilized two CIRS phantoms: an older specimen with increased hardness due to aging, and a newer phantom with softer material properties. Five anonymized CT scans from trauma patients (four male, one female) were used as a reference for phantom surface topology creation. These were full-body CT scans, with corresponding surface meshes for the skin, skeleton, and vasculature. The patients had varying BMI, as increased adipose tissue frequently complicates femoral arterial access. Clinically, vascular access targets the common femoral artery segment between the bifurcation and inguinal ligament. Unlike commercial phantoms with parallel vessels, human femoral arteries exhibit natural curvature and angulation. To enhance anatomical fidelity, we use the Meshmixer CAD software to precisely align the phantom’s bifurcated vessel with patient-specific anatomy by matching the 3D position of the bifurcation point and the orientation of the common femoral artery (see Figure 2). While each patient provided bilateral data, symmetry was assumed, and one side per patient was randomly selected for phantom creation.
Following alignment of the phantom’s bifurcated vessel with patient femoral artery data, the phantom volume is subtracted from the patient’s body volume (Figure 3). The resulting complementary volume is used to fabricate customized 10% w/w polyvinyl alcohol (PVA) hydrogel layers that replicate patient-specific surface topology when positioned atop the commercial phantoms. Fabrication involves a three-day freeze/one-day thaw cycle before mold removal. To achieve anatomically accurate positioning representative of supine patients, support platforms are designed based on patients CT data (Figure 3) that both orient the commercial phantom correctly and ensure precise positioning of the hydrogel layer relative to the CIRS models. Figure 4 compares the designed phantom models to patient anatomies using two metrics: the shortest $L_{2}$ distance from skin surface to bifurcation point, and depth ranges for both the acceptable arterial entry zone and complete artery. While patient arteries demonstrate greater depth variation due to natural curvature compared to the phantom’s planar vessel, both exhibit comparable depth ranges and significant variability by design (12-56mm) across all cases.
@@ III-B Femoral Artery Reconstruction
Accurate 3D reconstruction of the femoral artery is essential for identifying critical anatomical landmarks, such as the bifurcation, which help estimating safe needle entry points. Accurate localization of an entry zone enhances needle insertion planning, thereby minimizing complications such as arterial trauma or suboptimal catheter positioning. This study focuses specifically on achieving high-fidelity 3D reconstruction of the femoral artery using the methodology outlined below.
III-B1 Robotic US System
@@
The robotic system, shown in Figure 1, comprises a UR10e (Universal Robots) robotic manipulator with an integrated force/torque sensor, a convex US probe (C3 HD, Clarius Mobile Health) mounted on the robot’s end effector, and a RealSense camera (Intel Corporation) attached to a tripod.
@@
Given the predefined anatomical placement of the femoral artery within each phantom, the system bypasses vascular localization and instead focuses on autonomous scanning from one end of the artery to the other. In a real clinical scenario, this information can be provided to the robotic system by a paramedic. In the presented pipeline, this information is extracted using the phantom’s point cloud and pre-operative CT scan. The phantom’s surface geometry is first captured using the RealSense camera, generating a point cloud representation. Using the known start and end points of the artery (from CT scan), a straight-line trajectory is computed and projected onto the phantom’s point cloud within a plane that is both perpendicular to the best-fit surface plane of the phantom and contains the trajectory line. The US probe’s orientation along the scanning path is determined by the normal vector of the best-fit plane. The resulting scanning trajectory is discretized at 1mm intervals and provided as waypoints to the robot’s controller for execution.
@@
During scanning, the robotic system employs an impedance control strategy implemented through Universal Robots’ built-in force/torque control framework. This algorithm actively regulates the end-effector’s applied force, positional accuracy, and orientation, ensuring stable and consistent probe-surface contact. The scanning procedure was conducted with a contact force of 6 N for the older CIRS phantom and 4 N for the newer model. Throughout the scanning process, the robot’s end-effector poses and corresponding US images are recorded in a rosbag file. Synchronization is maintained using ROS2 timestamps to ensure precise temporal alignment between the positional data and acquired US frames.
@@
To ensure seamless integration of all system components, a unified reference frame is established, as illustrated in Figure 1. The notation $T^{A}_{B}$ denotes the homogeneous transformation that maps coordinates from frame $A$ to frame $B$. The essential reference frames employed in this system include: $T^{P}_{EE}$, the transformation describing the position and orientation of the US probe relative to the robot’s end effector; $T^{C}_{RB}$, the transformation between the camera frame and the robot’s base; $T^{US}_{EE}$, the transformation between the US image and the robot’s end effector; $T^{PH}_{CT}$, the transformation aligning the physical phantom with its corresponding CAD model reconstructed from computed tomography (CT) scans.
@@
US Image Calibration: $T^{US}_{EE}$ is determined through a point-based calibration method [33]. The procedure involves acquiring US images of a spatially fixed point in a water bath from multiple robot-configured poses, solving the constraint $B_{i}X^{-1}p_{i}=B_{j}X^{-1}p_{j}$. Calibration was performed at an US imaging depth of 100mm, maintained across all experiments. It is important to note that separate calibration matrices would be required for different imaging depths due to depth-dependent scaling variations.
@@
Eye-to-Hand Camera Calibration: A checkerboard was rigidly mounted to the robot table, and a calibrated tool tip guided to the checkerboard’s intersection points. The tool’s position relative to the robot base is recorded while the 3D camera captures depth images, extracting corresponding point cloud data. The transformation between the camera and robot base is computed using the Least Squares method.
@@
Phantom-to-Patient Registration: Scan start and end points are defined in patient-specific CAD models derived from the CT scans. To transfer this information to physical phantoms, an Iterative Closest Point (ICP) algorithm is used to compute the transformation between the phantom in the camera frame and its corresponding CAD model. The initial alignment is achieved by matching centroids and principal components before refining the registration with ICP.
III-B2 Vascular Segmentation NetworkVideos

### v2

@@ Abstract
[…]
@@ I Introduction
Femoral arterial access is integral to interventional medicine, enabling procedures such as diagnostic angiography, therapeutic catheterization, and emergency interventions. Its clinical utilization has grown substantially, with use of femoral access for transcatheter aortic valve replacement in the U.S. increasing from 57.1% in 2011 to 95.3% in 2019 [1]. Despite its broad adoption, femoral access remains among the most technically challenging vascular entry sites [2], with a 5.9% access failure rate reported in percutaneous aortic procedures [3]. The procedure’s complexity is compounded by anatomical variability, overlying adipose tissue, and the required accuracy in US image interpretation. Improper needle placement increases the risk of serious complications, including retroperitoneal hemorrhage, thrombosis, and other vascular sequelae [2, 4]. Femoral arterial access is hence currently performed only by skilled physicians in controlled hospital environments.
@@
Robotic systems can address aforementioned challenges by leveraging their precision and repeatability in US-guided vascular interventions. These systems can enable accurate 3D reconstruction of vasculature through automated US image segmentation, serving either as fully autonomous platforms, or as clinician-assistive technologies. Furthermore, they hold significant potential for deployment in resource-constrained environments where specialized healthcare personnel are unavailable, such as in combat zones, remote locations, or an ambulance. Although robotic US scanning and reconstruction is a mature research topic, most investigations have employed overly simplified phantom models that cannot represent the anatomical complexity encountered in clinical scenarios [5, 6], or relied on single custom-designed phantoms for both system development and validation [7, 8]. To the best of the authors’ knowledge, no prior study has validated robotic femoral artery reconstruction against an independent 3D ground truth [9, 10], a critical step for reliable vascular needle placement where full arterial geometry, including bifurcation and safe entry zones, is essential for the procedure’s success and patient’s safety.
In this manuscript, we present an autonomous robotic system for femoral artery scanning. The proposed system integrates US image acquisition, arterial segmentation, and 3D centerline reconstruction. In particular, this work’s main contributions are as follows: (1) demonstration of autonomous robotic scanning and 3D reconstruction of bifurcated femoral artery on five vascular phantoms created from real patient CT data; (2) development of a video-based segmentation network for axial artery segmentation in US video sequences; (3) comprehensive evaluation of the reconstructed bifurcated artery against ground-truth CT data. This study is the first to validate an autonomous robotic system for US scanning of the femoral artery on a diverse set of patient-specific phantoms, introducing a more rigorous framework for evaluating robotic performance in vascular imaging and intervention.
@@ II-1 Vascular Phantoms
Although commercial vascular phantoms provide good US image quality, their oversimplified geometries - such as rectangular prisms (CIRS model 072) or semi-cylinders (CAE Vascular Access Phantom) - fail to replicate realistic anatomy, hindering the transfer of robotic solutions or other approaches to real clinical procedures. Larger models such as the Blue Phantom Femoral Access trainer offer improved anatomical representation, but they are often prohibitively expensive. Perhaps most significantly, these phantoms typically derive from a single anatomical model, limiting their representation of human variation, particularly for individuals with higher body mass indices (BMI). Recently, wearable training systems combined with US simulation software, such as the SonoSkin trainer (Simulab, Seattle, WA), garnered more attention. The systems consist of a skin-like cover that can be placed over a mannequin or person, paired with a simulated US probe that triggers predetermined US images at specific anatomical locations. While such systems enable basic examination practice, their reliance on pre-loaded images and lack of orientation-specific feedback limits their utility in developing image acquisition skills, or advancing US image interpretation. Additionally, creating a custom vascular phantom requires self-healing materials and durable vessels to withstand repeated needle punctures, a technology largely protected as proprietary by commercial manufacturers. We provide a simple cost-effective solution to overcome these challenges. We develop five novel patient-specific phantoms by modifying the CIRS model 072 (Sun Nuclear, U.S.) vascular phantom’s surface topology and overall orientation to better represent realistic anatomy, while retaining the commercial phantom’s established acoustic and mechanical properties.
II-2 Ultrasound Vascular Segmentation
Automatic segmentation of US images acquired during robotic scanning presents significant challenges, particularly when integrated with force-control strategies. Although force control enhances safety during probe manipulation, it introduces variability in probe coupling and contact pressure. These effects, compounded by the inherently low signal-to-noise ratio and susceptibility to acoustic artifacts in US imaging, hinder accurate vascular segmentation.
Earlier works in US vessel segmentation used geometric and dynamic models, typically requiring seed points initialization in the first frame [11, 12, 13]. Other studies have improved performance by combining Doppler and B-mode US [14, 15, 16]. With the emergence of deep learning, most vessel segmentation approaches employed U-Net–based architectures that operate on individual US frames [17, 18, 19]. More recently, the introduction of transformers for modeling long-range dependencies has motivated the integration of attention mechanisms into U-Net–like models to incorporate temporal memory from US video sequences. However, to the best of author’s knowledge, most studies have focused on obstetrics [20], breast sonography [21], and echocardiography [22], with comparatively little focus on video-based vascular US segmentation in robotic scanning.
@@ II-3 Robotic Ultrasound Vascular Scanning
Robotic US has emerged as a promising solution to address several drawbacks of hand-held US imaging for abdominal [23, 24, 25, 26] and cardiac [27] procedures. In particular to vascular imaging, prior research has focused on vessel segmentation, tracking, and localization. Jiang et al. [28] used segmented vascular structures to align a robotic probe perpendicular to the anatomy and later refined spatial positioning using 3D point cloud data of the scanning surface [29]. Chen et al. [30] used color Doppler with B-mode images and analyzed them through deep learning for vascular access. Merouche et al. [7] conducted an automatic tracking of artery and 3D reconstruction by initializing the cross-sectional view of the artery. Ning et al. [31] employed a weakly supervised US vessel segmentation network for probe orientation estimation, and a force-guided reinforcement learning agent to maintain optimal probe angle. These systems, however, were validated on either commercial or custom-designed phantoms, which fail to fully capture the anatomical complexity of real humans. While some studies have validated their approaches on animals [9] or human subjects [31], they were primarily limited to US image acquisition, with no independent ground truth to assess 3D reconstruction quality. In contrast, our proposed robotic US vascular scanning system is validated on patient-specific phantoms with their corresponding CT scans, providing a more representative assessment of reconstruction quality.
This manuscript is structured as follows: Section III.A describes the methodology and rationale for vascular phantom development. Section III.B details the autonomous robotic US scanning protocol and introduces a novel video-based US segmentation network. Section IV presents and analyzes the experimental results from robotic US image acquisitions. Section V concludes the paper and outlines directions for future work.
@@ III-A Patient-Specific Vascular Phantoms
The phantoms comprise a commercial model with embedded vasculature as the base, overlaid with our custom-made hydrogel layers to more accurately replicate human anatomical surface topology. The commercial CIRS 072 model was used as the base phantom. It contains four embedded vessel structures: two lateral vessels of different diameters, and a central bifurcating vessel serving as our experimental target. To reduce detection artifacts, the blood-mimicking fluid was drained from the two lateral vessels. We utilized two CIRS phantoms: an older specimen whose hardness increased due to aging, and a newer phantom with softer material properties. Five anonymized trauma patient CT scans (four male, one female) served as references for generating phantom surface topology, with corresponding surface meshes of skin, skeleton, and vasculature. The patients had varying BMI, as increased adipose tissue commonly complicates femoral arterial access. Clinically, access targets the common femoral artery segment between the bifurcation and inguinal ligament. Unlike commercial phantoms with parallel vessels, human femoral arteries exhibit natural curvature. To enhance anatomical fidelity, we use the Meshmixer CAD software to precisely align the phantom’s bifurcated vessel with patient-specific anatomy by matching the 3D position of the bifurcation point and direction of the common femoral artery (see Figure 2). Although bilateral data were available for each patient, anatomical symmetry was assumed, and a single side was randomly selected for phantom generation.
Following alignment of the phantom’s bifurcating vessel with the patients’ femoral artery, the phantom volume is digitally subtracted from the patient’s body volume (see Figure 3). The resulting complementary volumes are used to create molds for customized 10% w/w polyvinyl alcohol hydrogel layers, which replicate patient-specific surface topology when placed atop the commercial phantoms. Hydrogel fabrication follows a three-day freeze/one-day thaw cycle, after which the hydrogel layer is demolded. To achieve anatomically accurate supine positioning, support platforms are designed from patient CT data (see Figure 3) to orient the commercial phantom, and ensure precise alignment of the hydrogel layer with the underlying CIRS models.
Figure 4 compares the designed phantom models to patient anatomies using two metrics: the shortest $L_{2}$ distance from skin surface to bifurcation point, and depth ranges for both the acceptable arterial entry zone and complete artery. While patient arteries demonstrate greater depth variation due to natural curvature compared to the phantom’s planar vessel, both exhibit comparable depth ranges and significant variability by design (12-56mm) across all cases.
@@ III-B Femoral Artery Reconstruction
Accurate 3D reconstruction of the femoral artery is essential for identifying critical landmarks, such as the bifurcation, that define safe needle entry points. Precise localization of the entry zone improves needle insertion planning and reduces complications, including arterial trauma and suboptimal catheter placement. This study focuses on achieving high-fidelity femoral artery reconstruction using the methodology described below.
III-B1 Robotic Ultrasound System
@@
The robotic system, shown in Figure 1, comprises a UR10e manipulator with an integrated force/torque sensor (Universal Robots, Denmark), a convex US probe (C3 HD, Clarius Mobile Health, Canada) mounted on the robot’s end effector, and a RealSense camera (Intel Corporation, U.S.) attached to a tripod.
@@
Given the predefined anatomical placement of the femoral artery within each phantom, the system bypasses vascular localization and instead focuses on autonomous scanning from one end of the artery to the other. In a real clinical scenario, this information can be provided to the robotic system by a paramedic. In the presented pipeline, this information is extracted using the phantom’s point cloud and pre-operative CT scan. The phantom’s surface is first scanned with a RealSense camera to produce a 3D point cloud. Using the known arterial start and end points from the CT scan, a straight-line trajectory is computed. The phantom surface is approximated by a best-fit plane, and a perpendicular plane passing through the straight-line trajectory is defined. The intersection of the perpendicular plane with the phantom surface point cloud defines the final US scanning trajectory, which is then discretized at 1mm intervals and supplied to the robot controller as execution waypoints. The US probe orientation is held constant along the scanning path, with its $z$-axis aligned to the normal of the best-fit plane (blue arrow on probe tip in Fig.1) and its $x$-axis (red arrow) aligned perpendicular to the initial straight-line trajectory, thereby fully defining the probe’s pose.
@@
During scanning, the robotic system employs a force-control strategy implemented through Universal Robots’ built-in force/torque control framework. This algorithm concurrently regulates the end-effector’s applied force, positional accuracy, and orientation. The scanning procedure was conducted with a contact force of 6 N for the older CIRS phantom, and 4 N for the newer model, consistent with values commonly reported in clinical practice for vascular imaging [32]. Throughout the scanning process, the robot’s end-effector poses and corresponding US images are recorded in a rosbag file. Synchronization is maintained using ROS2 timestamps to ensure temporal alignment between positional data and acquired US frames.
@@
To ensure seamless integration of all system components, a unified reference frame is established, as illustrated in Figure 1. The notation $T^{A}_{B}$ denotes the homogeneous transformation that maps coordinates from frame $A$ to frame $B$. The essential reference frames employed in this system include: $T^{P}_{EE}$, the transformation describing the position and orientation of the US probe relative to the robot’s end effector; $T^{C}_{RB}$, the transformation between the camera frame and the robot’s base; $T^{US}_{EE}$, the transformation between the US image and the robot’s end effector; $T^{PH}_{CT}$, the transformation aligning the physical phantom with its corresponding CAD model obtained from CT scans.
@@
US Image Calibration: $T^{US}_{EE}$ is determined through a point-based calibration method [33]. The procedure involves acquiring US images of a spatially fixed point in a water bath from multiple robot poses, and solving $B_{i}X^{-1}p_{i}=B_{j}X^{-1}p_{j}$ using least squares method. Calibration was performed at an US imaging depth of 100mm, maintained across all experiments. It is important to note that separate calibration matrices would be required for different imaging depths due to depth-dependent scaling variations.
@@
Eye-to-Hand Camera Calibration: A checkerboard was rigidly fixed to the robot table, and a calibration tool tip mounted on the robot end effector was guided to the checkerboard intersection points. The tool’s position relative to the robot base was recorded while the 3D camera captured a depth image of the checkerboard. The transformation between the camera and robot base was then estimated via least-squares.
@@
Phantom-to-Patient Registration: Scan start and end points are first defined on patient-specific CAD models reconstructed from CT scans. To map these points to the physical phantom, the transformation between the phantom in the camera frame and its corresponding CAD model is estimated. An initial alignment is obtained by matching centroids and principal component axes, followed by refinement using the Iterative Closest Point (ICP) algorithm.
III-B2 Vascular Segmentation Network